If C1=4.3 ?F, C2=5.1 ?F, C3=3.7 ?F and C4=6.5 ?F, what is the equivalent capacitance of this capacitor network?

If C1=4.3 ?F, C2=5.1 ?F, C3=3.7 ?F
Answer #1

C2 & C3 are in parallel so their equivalent capacitance = 5.1 + 3.7 = 8.8 \muo \muF

These are in series with C1 and C4 so the circuit's equivalent (total) capacitance = Ceq  

1/Ceq = 1/4.3 + 1/8.8 + 1/6.5 = 0.500040657

solving this, we get

Ceq = 1.999837 = 2 \muo \muF (approx)

so, the equivalent capacitance of this capacitor network is 2 \muo \muF